OpenStudy (anonymous):

what is the slope of the line that passes through the pair of points (3,-4) and (5,-7) a. 2/3 b. -2/3 c. -3/2 d. 3/2

OpenStudy (anonymous):

(-7-(-4))/(5-3)=(-7+4)/2=-3/2 option c

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

Slope of any line parallel to x axis is 0

OpenStudy (anonymous):

@juliannah337665